# The Dynamical History of the Kepler-221 Planet System

### v1

@@ 6.4 Application to other systems
Other exo-planetary systems with one planet outside the 1st-order resonance chain may also be described with ideas introduced in this work. In particular, we highlight the K2-138 and Kepler-402 systems with the information of these systems listed in Table 5 and Table 6. K2-138 is a K1-type star with six super-Earths. The period ratios between the inner five planets are close to 3:2 and all planets in K2-138 are in three-body resonance, with the innermost three pairs in $(p,q,r)=(2,3,5)$ 3BR. The outermost three planets (planets e, f, and g) are locked in a peculiar 1st-order $(2,3,4)$ 3BR, which is the first pure 1st-order 3BR detected in a multi-planetary system (Cerioni & Beaugé, 2023). Kepler-402 is an F2-type star with four super-Earths (planets b, c, d, and e). Similar to Kepler-221, planets b, c, and e are in a $(7,8,15)$ 3BR, with the inner period ratio close to 3:2 and the outer period ratio close to 16:9. Planet d is not in resonance with any other planet.
@@
The 3BRs in the K2-138 system are listed in Table 5. The inner three 3BRs pairs are close to $(2,3,5)$ 3BR, with the period ratio increasing from inner pairs to outer pairs. Therefore, it is reasonable to assume that the inner five planets are originally locked into 3:2 two-body resonances with each adjacent triple also locked in 3BRs. At 1.544 and 3.290 the period ratios of planets e, f, and g are relatively far from integer period ratios, which could be the result of orbital expansion along $(2,3,4)$ 1st-order 3BR from an inner 3:2 and outer 3:1 two-body resonance. The formation of 1st-order 3BRs was also seen in our simulations for the Kepler-221 system when we attempted to reform the b/c/e 3BR in the post-collision phase (see Section 4 and Appendix A). This demonstrates that the formation of pure 1st-order 3BR is possible in systems similar to Kepler-221. Compared to 0th-order 3BRs, a 1st-order 3BR does not originate from two two-body resonances and could form relatively far from exact 2-body commensurabilities (Petit, 2021) (see Appendix A). A possible formation scenario of K2-138 could be that the inner five planets are first locked in a chain of 3:2 two-body resonances. Then tidal damping expands the period ratio of the inner planet, and the outer planet g coincidentally comes close to the $(2,3,4)$ 1st-order 3BR and becomes trapped, similar to the resonance reformation process discussed in Section 4.3 and Appendix A.
@@
The period ratios of the Kepler-402 planetary system are shown in Table 6 with the corresponding closest 3BR. It is clear from the figure that the adjacent planets are far from 3BR with larger normalized B-value but planets b, c, and e are in the $(7,8,15)$ pure 3BR relatively far from integer period ratio, similar to the Kepler-221 system. This 3BR between planets b, c, and e might expand from an inner 3:2 and outer 16:9 two-body resonance. However, it is unlikely that such a high-order resonance can be formed with simple convergent migration in the disk phase. Also, the middle planet d makes the resonance formation between c and e harder (see Section 4). Therefore, the $(7,8,15)$ 3BR for planets b, c, and e probably does not originate from two-body resonance between c and e. Possibly, the Kepler-402 system formed similarly to Kepler-221 in the sense that that there were originally two planets between planets c and e ($\mathrm{d}_{1}$ and $\mathrm{d}_{2}$) with all five planets in a 1st-order resonance chain. The resonance number for planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ two-body resonance is very high (7:6 or 8:7 resonance, facilitating the resonances to break quickly after disk dispersal (Matsumoto et al., 2012), and resulting in the merging of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ into planet d. Afterward, tidal damping reforms the 3BR between planets b, c, and e, and the system period ratios expand to the observed value.
@@ 7 Conclusion
We have proposed a multi-phase formation model for the dynamical history of the Kepler-221 system, in line with its present architecture, of which the most peculiar feature is the out-of-resonance intermediate planet d. The envisioned scenario relies on two Ansatzes. First, there were originally five planets (planets b, c, $\mathrm{d}_{1}$, $\mathrm{d}_{2}$, and e) in a chain of 1st-order resonances. Second, after disk dispersal, the system experienced an instability, which broke all resonances and caused planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ to merge into planet d with its position consistent with its observed location. Under dissipative processes like tidal damping, the 6:3:1 three-body resonance between planets b, c, and e reformed. On evolutionary timescales, the period ratio of these three planets expanded under tidal dissipation until they reached the observed period ratio.
@@
    Immediately after the collision of d1 and d2, convergent migration between planets b and c is essential to reform the b/c/e 3BR. This implies a positive torque operating on planet b or an anomalously small tidal-Q of planet c. Stronger damping on planet c (possibly due to strong obliquity tides (Goldberg & Batygin, 2021)) would migrate planet c inward while a positive torque on planet b would migrate planet b outward. If the appropriate conditions are satisfied, the success rate amounts to 17% (see Figure 4).
@@
    If these conditions are not present, resonance reformation would fail or the b/c/e planets would end up in a $(1,3,3)$ 1st-order three-body resonance.
@@
    The properties of planet d are crucial for the outcome of the orbital expansion of the b/c/e sub-system. In order for the planets’ period ratios to evolve towards their current values, planets c, d, and e must have avoided the $(3,5,8)$ resonance, which would break the b, c, and e 3BR with a probability of 95%.
@@
    Therefore, planets b/c/e should have evolved along a line of slope $a_{\mathrm{cde}}>1.56$ in the period ratio plane of $P_{\mathrm{e}}/P_{\mathrm{d}}$ vs $P_{\mathrm{d}}/P_{\mathrm{c}}$. Conservation of angular momentum and the b/c/e 3BR angle gives an analytical expression for $a_{\mathrm{cde}}$ (see Equation (10)), which only depends on the mass ratios of planets b, c, and e.
@@
    The dynamical model proposed for the Kepler-221 system can be applied to other planetary systems in resonance. In particular, K2-138, which features three planets in a pure 1st-order 3BR, and Kepler-402, which, like Kepler-221, features an intermediate planet not part of the resonance chain.
@@ Acknowledgements.
We thank the referee for their constructive comments, which significantly improved the presentation of this work. This work is supported by the National Science Foundation of China under grant No. 12250610189 and 12233004. The authors gratefully thank Sharon Xuesong Wang, Fei Dai, Haochang Jiang, Yu Wang and Helong Huang for their useful discussions. Software: matplotlib (Hunter, 2007; Caswell et al., 2021), REBOUND (Rein & Liu, 2012), REBOUNDx (Tamayo et al., 2020).
Appendix A Trapping in first order 3BR
In the post-collision phase, planets b, c, and e can escape from 0th-order 3BR and instead get trapped in 1st-order 3BR, as shown in panels a and d in Figure 3. Although both originate from b/c 2:1 and c/e 3:1 2BR, the definitions of 0th-order and 1st-order 3BR for planets b, c, and e are different. The expression for b/c 2:1 outer and c/e 3:1 inner 2BR angle are:
@@ Appendix A Trapping in first order 3BR
$\phi_{\mathrm{ce,i}}=\lambda_{\mathrm{c}}-3\lambda_{\mathrm{e}}+2\varpi_{\mathrm{c}}$ (17)
@@
where $\lambda_{\mathrm{i}}$ is for mean longitude for the planet $i$ and $\varpi_{\mathrm{c}}$ represents the longitude of periapsis of planet c. For the definition of 0th-order 3BR, $\varpi_{c}$ is eliminated:
$\phi_{\mathrm{bce,0}}=2\phi_{\mathrm{bc}}-\phi_{\mathrm{ce}}=2\lambda_{\mathrm{b}}-5\lambda_{\mathrm{c}}+3\lambda_{\mathrm{e}}$ (18)
@@
The definition of a 1st-order 3BR contains the argument of periapsis of one planet and requires that $p+q-r=1$ (see Equation (2)). One example of 1st-order 3BR for planets b, c, and e is:
$\phi_{\mathrm{bce,1}}=\phi_{\mathrm{bc}}-\phi_{\mathrm{ce}}=\lambda_{\mathrm{b}}-3\lambda_{\mathrm{c}}+3\lambda_{\mathrm{e}}-\varpi_{\mathrm{c}}$ (19)
@@
If planet b/c and c/e are in a 2BR resonance (i.e., $\phi_{\mathrm{bc,o}}$ and $\phi_{\mathrm{ce,i}}$ librate), both $\phi_{\mathrm{bce,0}}$ and $\phi_{\mathrm{bce,1}}$ librate because these two 3BR angles are linear combination of the 2BR angle. In this case, planets b, c, and e are trapped in $(2,3,5)$ 0th-order 3BR (see Equation (2)). If $\phi_{\mathrm{bce,1}}$ liberates, but not the 0th-order 3BR angle and 2BR angles, then planets b, c, and e are trapped in $(1,3,3)$ 1st-order 3BR. The strength of 0th and 1st-order 3BRs are roughly determined by $p+q$, with the smaller value representing stronger resonance (Petit 2021).
@@
Figure 10: Period ratios and resonance angles for planets b, c, and e trapped in 0th and 1st-order 3BR in post-collision phase. Panels a and b represent the period ratio evolution with the color bar indicating time. The black and red dashed lines respectively correspond to $(2,3,5)$ 0th-order 3BR and $(1,3,3)$ 1st-order 3BR following Equation (7). Panel c and d plot the evolution of the 0th-order 3BR angle $\phi_{\mathrm{bce,0}}=2\lambda_{\mathrm{b}}-5\lambda_{\mathrm{c}}+3\lambda_{\mathrm{e}}$ and 1st-order 3BR angle $\phi_{\mathrm{bce,1}}=\lambda_{\mathrm{b}}-3\lambda_{\mathrm{c}}+3\lambda_{\mathrm{e}}-\varpi_{\mathrm{c}}$. Panel e and f plot the b/c outer resonance angle $\phi_{\mathrm{bc,o}}=\lambda_{\mathrm{b}}-2\lambda_{\mathrm{c}}+\varpi_{\mathrm{c}}$ and c/e inner resonance angle $\phi_{\mathrm{ce,i}}=\lambda_{\mathrm{c}}-3\lambda_{\mathrm{e}}+2\varpi_{\mathrm{c}}$. The right panels correspond to the simulation where a positive torque is applied on planet b and planets b, c, and e get trapped in 0th-order 3BR. In contrast, in the left panels, there is no torque on planet b and planets are trapped in a 1st-order 3BR.
@@
Generally, the formation of 0th-order 3BR is an inevitable result of 2BR formation, so planets should first be captured when their period ratios are close to integer ratios. If planets are far away from integer ratio, then they are more likely to be captured into the stronger 1st-order 3BR with smaller $p+q$ values for planets b, c, and e. Figure 10 illustrates the process of resonance formation in the post-collision phase (see Section 4). The left panels show the simulation where no additional torque is applied on planet b and the tidal $Q$ parameter is the same for all planets. Planet b, c, and e skip the 0th-order 3BR and get trapped into 1st-order 3BR. This is because planets b and c did not first reform 2BR because they do not undergo convergent migration (Petit 2021). So the planets period ratio expands skipping the 0th-order resonance (black dashed line in panel a), and forms the stronger 1st-order 3BR along the red dashed line. Therefore, only the 1st-order resonance angle librates (see panels c and d in Figure 10).
@@
The right panel in Figure 10 corresponds to a simulation with an additional positive torque on planet b (see Section 4.3), which results in the outward migration of planet b. This decreases the period ratio between planets b and c (panel b) and leads to convergent migration between these two planets, and first reform b/c 2BR near integer ratio around 1.5 Myr (panel f). This promotes the reformation of the 0th-order 3BR, which forms simultaneously with the c/e 2BR (panel d and f) around 1.7 Myr.
@@
The simulations show that the formation of pure 1st-order 3BR is possible if planets are close to strong 1st-order 3BR position due to resonance breaking event. Similar scenarios could be applied to the K2-138 system, which hosts the only pure 1st-order 3BR resonance identified in the exoplanet census (Cerioni & Beaugé 2023). In the K2-138 system, the outermost three planets are trapped in a pure 1st-order 3BR. Possibly, the formation of K2-138 1st-order 3BR is similar to the resonance formation process in Kepler-221, as shown in Figure 10 (see Section 6.4).
@@ Appendix B Optimized Collision Phase
Following the model introduced in Section 2, we proceed with simulation in chronological order with the optimized model, starting from the collision phase. Panels a and c in Figure 11 show an example simulation of the optimized collision phase. We initialize the five planets in 1st-order resonance with masses according to Model 1 in Table 2 (Chen & Kipping 2017). In panel a, the b/c/e 3BR angle is first on a fixed value different from 0 or $\pi$ because multiple 3BR are present for the five planets in the system. In the optimized collision phase, planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ are removed by decreasing their masses adiabatically with a timescale of 1 Myr until their masses reach zero. When $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ are removed, only b/c/e 3BR is left and it goes to the isolated fixed point librating around $\pi$ (Delisle 2017). After this, planet d is inserted at its current position (panel a) and its mass adiabatically increases until $m_{\mathrm{d}}$ reaches the value shown in Table 2 with a timescale of 0.3 Myr, which ensures that the 3BR between planets b, c, and e do not break due to the mass change of planet d.

### v2

@@ 6.4 Application to other systems
Other exo-planetary systems with one planet outside the first-order resonance chain may also be described with ideas introduced in this work. In particular, we highlight the K2-138 and Kepler-402 systems with the information of these systems listed in Table 5 and Table 6. K2-138 is a K1-type star with six super-Earths. The period ratios between the inner five planets are close to 3:2 and all planets in K2-138 are in three-body resonance, with the innermost three pairs in $(p,q,r)=(2,3,5)$ 3BR. The outermost three planets (planets e, f, and g) are locked in a peculiar first-order $(2,3,4)$ 3BR, which is the first pure first-order 3BR detected in a multi-planetary system (Cerioni & Beaugé, 2023). Kepler-402 is an F2-type star with four super-Earths (planets b, c, d, and e). Similar to Kepler-221, planets b, c, and e are in a $(7,8,15)$ 3BR, with the inner period ratio close to 3:2 and the outer period ratio close to 16:9. Planet d is not in resonance with any other planet.
@@
The 3BRs in the K2-138 system are listed in Table 5. The inner three 3BRs pairs are close to $(2,3,5)$ 3BR, with the period ratio increasing from inner pairs to outer pairs. Therefore, it is reasonable to assume that the inner five planets are originally locked into 3:2 two-body resonances with each adjacent triple also locked in 3BRs. At 1.544 and 3.290 the period ratios of planets e, f, and g are relatively far from integer period ratios, which could be the result of orbital expansion along $(2,3,4)$ first-order 3BR from an inner 3:2 and outer 3:1 two-body resonance. The formation of first-order 3BRs was also seen in our simulations for the Kepler-221 system when we attempted to reform the (b, c, e) 3BR in the post-collision phase (see Section 4 and Appendix A). This demonstrates that the formation of pure first-order 3BR is possible in systems similar to Kepler-221. Compared to zeroth-order 3BRs, a first-order 3BR does not originate from two two-body resonances and could form relatively far from exact 2-body commensurabilities (Petit, 2021) (see Appendix A). A possible formation scenario of the K2-138 could be that the inner five planets are first locked in a chain of 3:2 two-body resonances. Then tidal damping expands the period ratio of the inner planet, and the outer planet g coincidentally comes close to the $(2,3,4)$ first-order 3BR and becomes trapped, similar to the resonance reformation process discussed in Section 4.3 and Appendix A.
@@
The period ratios of the Kepler-402 planetary system are shown in Table 6 with the corresponding closest 3BR. It is clear from the figure that the adjacent planets are far from 3BR with larger normalized B-value but planets b, c, and e are in the $(7,8,15)$ pure 3BR relatively far from integer period ratio, similar to the Kepler-221 system. This 3BR between planets b, c, and e might expand from an inner 3:2 and outer 16:9 two-body resonance. However, it is unlikely that such a high-order resonance can be formed with simple convergent migration in the disk phase. Also, the middle planet d makes the resonance formation between c and e harder (see Section 4). Therefore, the $(7,8,15)$ 3BR for planets b, c, and e probably does not originate from two-body resonance between c and e. Possibly, the Kepler-402 system formed similarly to Kepler-221 in the sense that that there were originally two planets between planets c and e ($\mathrm{d}_{1}$ and $\mathrm{d}_{2}$) with all five planets in a first-order resonance chain. The resonance number for planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ two-body resonance is very high (7:6 or 8:7 resonance, facilitating the resonances to break quickly after disk dispersal (Matsumoto et al., 2012), and resulting in the merging of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ into planet d. Afterward, tidal damping reforms the 3BR between planets b, c, and e, and the system period ratios expand to the observed value.
@@ 7 Conclusion
We have proposed a multi-phase formation model for the dynamical history of the Kepler-221 system, in line with its present architecture, of which the most peculiar feature is the out-of-resonance intermediate planet d. The envisioned scenario relies on two Ansatzes. First, there were originally five planets (planets b, c, $\mathrm{d}_{1}$, $\mathrm{d}_{2}$, and e) in a chain of first-order resonances. Second, after disk dispersal, the system experienced an instability, which broke all resonances and caused planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ to merge into planet d with its position consistent with its observed location. Under dissipative processes such as tidal damping, the 6:3:1 three-body resonance between planets b, c, and e reformed. On evolutionary timescales, the period ratio of these three planets expanded under tidal dissipation until they reached the observed period ratio.
@@
    Immediately after the collision of d1 and d2, convergent migration between planets b and c is essential to reform the (b, c, e) 3BR. This implies a positive torque operating on planet b or an anomalously small tidal-Q of planet c. Stronger damping on planet c (possibly due to strong obliquity tides (Goldberg & Batygin, 2021)) would migrate planet c inward while a positive torque on planet b would migrate planet b outward. If the appropriate conditions are satisfied, the success rate amounts to 17% (see Figure 4).
@@
    If these conditions are not present, resonance reformation would fail or the (b, c, e) planets would end up in a $(1,3,3)$ first-order three-body resonance.
@@
    The properties of planet d are crucial for the outcome of the orbital expansion of the (b, c, e) sub-system. In order for the planets’ period ratios to evolve toward their current values, planets c, d, and e must have avoided the $(3,5,8)$ resonance, which would break the b, c, and e 3BR with a probability of 95%.
@@
    Therefore, planets (b, c, e) should have evolved along a line of slope $a_{\mathrm{cde}}>1.56$ in the period ratio plane of $P_{\mathrm{e}}/P_{\mathrm{d}}$ vs $P_{\mathrm{d}}/P_{\mathrm{c}}$. Conservation of angular momentum and the (b, c, e) 3BR angle gives an analytical expression for $a_{\mathrm{cde}}$ (see Equation (10)), which only depends on the mass ratios of planets b, c, and e.
@@
    The dynamical model proposed for the Kepler-221 system can be applied to other planetary systems in resonance, in particular, K2-138, which features three planets in a pure first-order 3BR, and Kepler-402, which, as with Kepler-221, features an intermediate planet not part of the resonance chain.
@@ Acknowledgements.
We thank the referee for their constructive comments, which significantly improved the presentation of this work. This work is supported by the National Natural Science Foundation of China under grant No. 12250610189, 12233004 and 12473065. The authors gratefully thank Sharon Xuesong Wang, Fei Dai, Haochang Jiang, Yu Wang and Helong Huang for their useful discussions. Software: matplotlib (Hunter, 2007; Caswell et al., 2021), REBOUND (Rein & Liu, 2012), REBOUNDx (Tamayo et al., 2020).
Appendix A Trapping in first-order 3BR
In the post-collision phase, planets b, c, and e can escape from zeroth-order 3BR and instead get trapped in first-order 3BR, as shown in panels a and d in Figure 3. Although both originate from b/c 2:1 and c/e 3:1 2BR, the definitions of zeroth-order and first-order 3BR for planets b, c, and e are different. The expression for b/c 2:1 outer and c/e 3:1 inner 2BR angle are:
@@ Appendix A Trapping in first-order 3BR
$\phi_{\mathrm{ce,i}}=\lambda_{\mathrm{c}}-3\lambda_{\mathrm{e}}+2\varpi_{\mathrm{c}},$ (17)
@@
where $\lambda_{\mathrm{i}}$ is for mean longitude for the planet $i$ and $\varpi_{\mathrm{c}}$ represents the longitude of periapsis of planet c. For the definition of zeroth-order 3BR, $\varpi_{c}$ is eliminated:
$\phi_{\mathrm{bce,0}}=2\phi_{\mathrm{bc}}-\phi_{\mathrm{ce}}=2\lambda_{\mathrm{b}}-5\lambda_{\mathrm{c}}+3\lambda_{\mathrm{e}},$ (18)
@@
The definition of a first-order 3BR contains the argument of periapsis of one planet and requires that $p+q-r=1$ (see Equation (2)). One example of first-order 3BR for planets b, c, and e is:
$\phi_{\mathrm{bce,1}}=\phi_{\mathrm{bc}}-\phi_{\mathrm{ce}}=\lambda_{\mathrm{b}}-3\lambda_{\mathrm{c}}+3\lambda_{\mathrm{e}}-\varpi_{\mathrm{c}},$ (19)
@@
If planet b/c and c/e are in a 2BR resonance (i.e., $\phi_{\mathrm{bc,o}}$ and $\phi_{\mathrm{ce,i}}$ librate), both $\phi_{\mathrm{bce,0}}$ and $\phi_{\mathrm{bce,1}}$ librate because these two 3BR angles are linear combination of the 2BR angle. In this case, planets b, c, and e are trapped in $(2,3,5)$ zeroth-order 3BR (see Equation (2)). If $\phi_{\mathrm{bce,1}}$ liberates, but not the zeroth-order 3BR angle and 2BR angles, then planets b, c, and e are trapped in $(1,3,3)$ first-order 3BR. The strength of 0th and first-order 3BRs are roughly determined by $p+q$, with the smaller value representing stronger resonance (Petit 2021).
@@
Figure 10: Period ratios and resonance angles for planets b, c, and e trapped in 0th and first-order 3BR in post-collision phase. Panels a and b represent the period ratio evolution with the color bar indicating time. The black and red dashed lines respectively correspond to $(2,3,5)$ zeroth-order 3BR and $(1,3,3)$ first-order 3BR following Equation (7). Panel c and d plot the evolution of the zeroth-order 3BR angle $\phi_{\mathrm{bce,0}}=2\lambda_{\mathrm{b}}-5\lambda_{\mathrm{c}}+3\lambda_{\mathrm{e}}$ and first-order 3BR angle $\phi_{\mathrm{bce,1}}=\lambda_{\mathrm{b}}-3\lambda_{\mathrm{c}}+3\lambda_{\mathrm{e}}-\varpi_{\mathrm{c}}$. Panel e and f plot the b/c outer resonance angle $\phi_{\mathrm{bc,o}}=\lambda_{\mathrm{b}}-2\lambda_{\mathrm{c}}+\varpi_{\mathrm{c}}$ and c/e inner resonance angle $\phi_{\mathrm{ce,i}}=\lambda_{\mathrm{c}}-3\lambda_{\mathrm{e}}+2\varpi_{\mathrm{c}}$. The right panels correspond to the simulation where a positive torque is applied on planet b and planets b, c, and e get trapped in zeroth-order 3BR. In contrast, in the left panels, there is no torque on planet b and planets are trapped in a first-order 3BR.
@@
Generally, the formation of zeroth-order 3BR is an inevitable result of 2BR formation, so planets should first be captured when their period ratios are close to integer ratios. If planets are far away from integer ratio, then they are more likely to be captured into the stronger first-order 3BR with smaller $p+q$ values for planets b, c, and e. Figure 10 illustrates the process of resonance formation in the post-collision phase (see Section 4). The left panels show the simulation where no additional torque is applied on planet b, and the tidal $Q$ parameter is the same for all planets. Planet b, c, and e skip the zeroth-order 3BR and get trapped into first-order 3BR. This is because planets b and c did not first reform 2BR because they do not undergo convergent migration (Petit 2021). So the planets period ratio expands skipping the zeroth-order resonance (black dashed line in panel a), and forms the stronger first-order 3BR along the red dashed line. Therefore, only the first-order resonance angle librates (see panels c and d in Figure 10).
@@
The right panel in Figure 10 corresponds to a simulation with an additional positive torque on planet b (see Section 4.3), which results in the outward migration of planet b. This decreases the period ratio between planets b and c (panel b) and leads to convergent migration between these two planets, and first reform b/c 2BR near integer ratio around 1.5 Myr (panel f). This promotes the reformation of the zeroth-order 3BR, which forms simultaneously with the c/e 2BR (panel d and f) around 1.7 Myr.
@@
The simulations show that the formation of pure first-order 3BR is possible if planets are close to strong first-order 3BR position due to resonance breaking event. Similar scenarios could be applied to the K2-138 system, which hosts the only pure first-order 3BR resonance identified in the exoplanet census (Cerioni & Beaugé 2023). In the K2-138 system, the outermost three planets are trapped in a pure first-order 3BR. Possibly, the formation of K2-138 first-order 3BR is similar to the resonance formation process in Kepler-221, as shown in Figure 10 (see Section 6.4).
@@ Appendix B Optimized collision phase
Following the model introduced in Section 2, we proceed with simulation in chronological order with the optimized model, starting from the collision phase. Panels a and c in Figure 11 show an example simulation of the optimized collision phase. We initialize the five planets in first-order resonance with masses according to Model 1 in Table 2 (Chen & Kipping 2017). In panel a, the (b, c, e) 3BR angle is first on a fixed value different from 0 or $\pi$ because multiple 3BR are present for the five planets in the system. In the optimized collision phase, planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ are removed by decreasing their masses adiabatically with a timescale of 1 Myr until their masses reach zero. When $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ are removed, only (b, c, e) 3BR is left and it goes to the isolated fixed point librating around $\pi$ (Delisle 2017). After this, planet d is inserted at its current position (panel a) and its mass adiabatically increases until $m_{\mathrm{d}}$ reaches the value shown in Table 2 with a timescale of 0.3 Myr, which ensures that the 3BR between planets b, c, and e do not break due to the mass change of planet d.